OpenStudy (31356):

(4x^2 y^3 z^4)^3 (-3xy) / (-2xy^2 z^3)^2 MEDAL REWARDED!! ;)

OpenStudy (31356):

@jim_thompson5910 @MelissaHolmes

OpenStudy (anonymous):

@jim_thompson5910 you got this... I got to go.... sorry

OpenStudy (31356):

Okay.

jimthompson5910 (jim_thompson5910):

can you simplify (4x^2 y^3 z^4)^3 for me?

jimthompson5910 (jim_thompson5910):

Or are you completely stuck here?

OpenStudy (31356):

I am completely stuck

jimthompson5910 (jim_thompson5910):

what is 4^3?

OpenStudy (31356):

Or else I wouldn't bother to answer the question myself.

OpenStudy (31356):

64

jimthompson5910 (jim_thompson5910):

ok, there's no need to be rude now

jimthompson5910 (jim_thompson5910):

now compute (x^2)^3 by using the rule (a^b)^c = a^(b*c)

OpenStudy (31356):

No, I think you misunderstood me, I was just saying it in a way, you know. t's polite in other countries you know.

jimthompson5910 (jim_thompson5910):

(x^2)^3 = ???

OpenStudy (31356):

Would the x stand for 1

OpenStudy (31356):

Or an understood of 1?

jimthompson5910 (jim_thompson5910):

it has a coefficient of 1, yes

OpenStudy (31356):

Okay

OpenStudy (31356):

8

jimthompson5910 (jim_thompson5910):

(x^2)^3 = x^(2*3) = x^6 using that rule

OpenStudy (31356):

Okay

jimthompson5910 (jim_thompson5910):

(y^3)^3 = ???

OpenStudy (31356):

3??????????

jimthompson5910 (jim_thompson5910):

use the rule posted above

jimthompson5910 (jim_thompson5910):

(a^b)^c = a^(b*c)

OpenStudy (31356):

Okay

jimthompson5910 (jim_thompson5910):

so (y^3)^3 = ???

OpenStudy (31356):

I understand the rest now, thanks, got to go.

jimthompson5910 (jim_thompson5910):

alright